OpenStudy (anonymous):

How do you solve -5(t-2)-(t+7)=4

OpenStudy (anonymous):

are you solving for t? if so, distribute the -5 to (t-2) and -1 to (t+7) and then add like terms t(-5)-2(-5)+t(-1)+7(-1)

OpenStudy (anonymous):

-5(t-2)-(t+7)=4 <-Problem -5t+10-t-7=4 <-Combine like terms -5t-t+10-7=4 -6t+3=4 <-subtract 3 -6t=1 <-Divide -6 to both sides t=-1/6 <-answer